### v2

# Mitigating Omitted Variable Bias in Empirical Software Engineering

Abstract.
Omitted variable bias occurs when a statistical model leaves out variables that are relevant determinants of the studied effects. This results in the model attributing the missing variables’ effect to some of the included variables—hence over- or under-estimating the latter’s true effect. Omitted variable bias presents a significant threat to the validity of empirical research, particularly in non-experimental studies such as those common in empirical software engineering.
This paper illustrates the impact of omitted variable bias on two case studies in the software engineering domain, and uses them to present methods to investigate the possible presence of omitted variable bias, to estimate its impact, and to mitigate its drawbacks. The analysis techniques we present are based on causal structural models of the variables of interest, which provide a practical, intuitive summary of the key relations among variables.
This paper demonstrates a sequence of analysis steps that inform the design and execution of similar empirical studies in software engineering. An important observation is that it pays off to invest effort investigating omitted variable bias before actually executing an empirical study, because this effort can lead to a more solid study design, and to a reduction in its threats to validity.
1. Introduction
In this day and era of big data analytics, where massive datasets are commonly available with scalable machine learning techniques to boot, it is tempting to believe that “more data” is a cure-all. Unfortunately, there are scenarios where training with more data cannot improve the quality of a statistical analysis—in fact, it may even worsen it.
Consider a common task of any empirical discipline: estimating the (average) effect that changing a variable $X$ has on another, dependent variable $Y$. In experimental study design parlance, $X$ is the treatment, and $Y$ is the outcome or response. Here is a concrete example in the domain of software engineering, which we will look into more closely in Section 3: estimating how using a different programming language (treatment $X$) affects the quality (outcome $Y$) of a program implemented in that language. If we have observational data about $X$ and $Y$, we may simply fit a statistical model—anything from a simple linear regression to fancier devices—on these data, and interpret the fitted model’s parameters connecting $X$ and $Y$ as an estimate of this effect. In this process, a well-known snag is omitted variable bias: if there exists another unmeasured variable $Z$ that affects both $X$ and $Y$, our estimate of the effect $X\to Y$ will spuriously include the combined effect of $X$ and $Z$ instead of the effect of $X$ alone. Continuing our empirical software engineering example, $Z$ may be a programmer’s intrinsic skills, which are all but sure to also affect the quality $Y$ of the written programs. In such a scenario, adding more data about $X$ and $Y$ (i.e., sampling more datapoints) is not going to help; in fact, it may just entrench our reliance on the biased estimate by reducing its variance and giving the false impression of reliability or “significance”.
Omitted variable bias is a widespread risk of any statistical analysis of observational data, regardless of whether one employs frequentist [22, 19, 47], Bayesian [26, 37, 14], or other kinds of machine learning models [1, 33]. In fact, it is always possible that some relevant variable was not measured, because it was unknown, inaccessible, or impractical, time-consuming, or expensive to measure with reasonable accuracy. This paper’s main contribution is presenting several mitigation strategies to cope with omitted variable bias, and demonstrating them in scenarios and examples that are relevant for software engineering empirical data analysis.
Of course, the ideal approach to avoid omitted variable bias is running a fully controlled experiment, where the treatment $X$ is assigned randomly, and the corresponding values of $Y$ are recorded. Controlled experiments are the gold standard in science precisely because they protect from omitted variable bias even when we don’t even know which other unmeasured variables may bias our estimate—since randomly assigning the treatment effectively removes its dependencies on any (unmeasured) confounder. The obvious reason why controlled experiments are not more common is that they are generally very expensive to run. In a field like empirical software engineering, proper controlled experiments are prohibitively challenging to design and run on the time scale of real-world software development—as opposed to “toy” programming tasks. In contrast, there is abundant observational data from software repositories that span large systems developed over years by many developers; but discovering genuine causal effects using these purely observational data must contend with omitted variable bias. Therefore, this paper focuses on mitigation strategies for omitted variable bias when analyzing observational data.
Omitted variable bias is a common and relevant problem—especially for empirical software data. However, it is by no means the only pitfall of analyzing empirical data: as we further discuss in Section 2, there are plenty of other challenges such as the included variable bias [15], “precisely inaccurate” analyses that hide biases [31], and unrepresentative population samples [5]. While each challenge requires different measures, they all likely involve trade-offs between costs and benefits—similarly to the present paper’s outlook on dealing with omitted variable bias.
Contributions.
This paper makes the following contributions:
  • Demonstrates the significance of omitted variable bias when analyzing empirical software engineering (observational) data;
  • Presents statistical methods to detect, quantify, and mitigate omitted variable bias when analyzing empirical data;
  • Provides simple guidelines for empirical software engineering researchers to apply those mitigation strategies in practical settings;
  • For reproducibility, all data and analysis scripts are available online:
    replication package:  https://figshare.com/s/fe607d8eb7c4cedbac75[16]
In a nutshell, the main aims of this paper are methodological: the case studies of Section 3 and Section 4 are based on real data, but are not meant to lead to novel findings about their domains. Instead, they demonstrate plausible scenarios where confounding may occur, and to explore different mitigation strategies.
Organization.
Section 2 first presents relevant related work on the topics of mitigating biases in statistical analysis and, more generally, best statistical practices for empirical software engineering; then, it introduces the key notations and concepts that the rest of the paper relies on. The rest of the paper demonstrates how to deal with omitted variable bias in two case studies: Section 3 uses a comparatively simpler model to estimate the effect of programming languages on code quality (also mentioned in the introduction), which serves as a relatively uncomplicated scenario. Then, Section 4 uses a more complex model to analyze the relation between team size (how many developers work on a project) and effort (how long it takes to complete the project). Although the two case studies differ in complexity, they are both based on realistic scenarios and models that we investigated in our previous work [14, 11]. Section 5 serves as a high-level summary of the whole article, presenting guidelines, in the form of a sequence of analysis steps that generate fundamental information to support the design of a study that mitigates omitted variable bias. Finally, Section 6 concludes the paper with a short summary of the main contributions.
2. Related Work and Background
This section starts (in Section 2.1) with a brief overview of the origins of causal analysis techniques for observational data, and how they have been adopted in empirical sciences (including software engineering). Then, Section 2.2 and Section 2.3 introduce the key concepts (causal relations, DAGs, confounders) of causal analysis that we will develop in the rest of the paper. Finally, Section 2.4 positions the paper’s contributions by relating them to other forms of confounding and causal inference bias.
2.1. Related Work
One of science’s ultimate goals is understanding the processes that underlie observed phenomena. This means discovering cause/effect relations between variables, as opposed to mere statistical associations. While the roots of causal analysis date back to Neyman’s potential outcomes framework [45], a comprehensive understanding of causality has emerged only later, in the late part of the 20th century. The key milestones in the development of a robust understanding of causality include: (1) Rubin [40]built upon Neyman’s pioneering work, introducing a framework for causality in nonrandomized (observational) studies. (2) Angrist et al. [2]introduced instrumental variables [^1] In a nutshell, an instrumental variable is a variable that acts like a natural experiment on the treatment. based on the Neyman-Rubin framework. (3) Working at about the same time as Rubin, Pearl started focusing on structural (graph) models [35], which culminated in his celebrated techniques for rigorously analyzing causal effects and confounding Pearl [36].
Among Pearl’s work, directed acyclic graphs (DAGs) have become widely used to model the structural dependencies between observed (and unobserved) variables. As we will demonstrate already with the simple example of Section 2.2, DAGs are, first of all, a practical notation to specify causal relations. They also support techniques to estimate the strength of the causal relation among some nodes in the graph. Usually this is done by constructing a (linear) statistical model among variables, selected according to the DAG’s structure. (We will demonstrate this shortly in Section 2.3.) Nowadays, causal analysis based on DAGs is routinely used in disciplines with a strong empirical component such as medicine [54, 46], epidemiology [18], economics [23], and biology [24].
A key question when working with DAGs is how to build a realistic DAG in the first place. In fields where the underlying fundamental mechanisms are well understood, a DAG can be built based on expert knowledge and previous work. Another approach is causal discovery (also called structural learning), which tries to identify causal relations from data [44]. While causal discovery algorithms have made significant progress in the last decade [56]—in part on the wave of the recent machine learning boom—they remain heuristic approaches that work correctly only under precise assumptions about the possible interactions. This limitation is intrinsic, as one of the fundamental results of Pearl’s framework is that causal relations cannot be inferred from data alone (at least not without limiting assumptions). Regardless of whether they come from expert knowledge, are inferred heuristically from data, or simply encode some (plausible) hypotheses, DAGs remain a practical tool to precisely denote, validate, and reason about the causal relations in a system.
While causal analysis techniques are not widely used in software engineering empirical research, they are gradually gaining traction. Siebert [43]’s recent survey reports 31 studies in empirical software engineering that targeted some kind of causal analysis technique. All of the reviewed papers were published in the last 15 years, which confirms that causal analysis is not yet an established practice but is slowly gaining popularity. The majority (17 out of 31) of papers reviewed by Siebert are about fault localization and refer to Baah et al. [3]—the first contribution that tried to apply a causal view instead of the traditional, purely correlational analysis that is commonplace in fault localization techniques [55, 57, 39]. Testing is another area targeted by several of the studies reviewed by Siebert; these include applications to mutation testing [25], simulation testing [7], and A/B testing [28]. The other papers reviewed by Siebert target various topics such as performance analysis (in one case still linked to fault localization [42]).
In the last few years, some more empirical software engineering research was published, targeting varied topics such as: (1) modeling rules of human knowledge and how they are made available to artificial intelligence systems [21]; (2) studying the impact of social media posts on the popularity of open-source projects [10]; (3) analyzing dependencies in configurable software systems [20]; (4) studying the impact of programming languages on coding competitions [15].
In conclusion, there is growing interest in understanding the concepts of causal analysis and applying them to analyzing software engineering data. The present paper further supports this trend by demonstrating how the framework of causal analysis can help mitigate the pervasive issue of omitted variable bias.
$X$ $\sim\textsf{Normal}(0,1)$
$\epsilon$ $\sim\textsf{Normal}(0,1)$
$Y$ $=bX+\epsilon$
XXYYZZ
(a) Process $p_{1}$: $Y$ depends on $X$, with some measurement error $\epsilon$.
$X$ $\sim\textsf{Normal}(0,1)$
$Z$ $\sim\textsf{Normal}(0,1)$
$\epsilon$ $\sim\textsf{Normal}(0,1)$
$Y$ $=bX+cZ+\epsilon$
XXYYZZ
(b) Process $p_{2}$: $Y$ depends on $X$ and $Z$, with some measurement error $\epsilon$.
$Z$ $\sim\textsf{Normal}(0,1)$
$X$ $=dZ+\epsilon_{X}$
$Y$ $=bX+cZ+\epsilon_{Y}$
$\epsilon_{X,Y}$ $\sim\textsf{Normal}(0,1)$
XXYYZZ
(c) Process $p_{3}$: $Y$ depends on $X$ and $Z$, and $X$ also depends on $Z$, with some measurement error $\epsilon$.
Figure 1. Three possible processes where variable $Y$ depends on variables $X$ and $Z$, and the corresponding DAGs capturing these structural relations.
$Y_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+\beta X_{i}$
(a) Model $m_{1}$: $Y$ is conditioned on $X$ only.
$Y_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+\beta X_{i}+\gamma Z_{i}$
(b) Model $m_{2}$: $Y$ is conditioned on both $X$ and $Z$.
Figure 2. Two linear regression models that capture the dependence between $X$, $Y$, and $Z$.
2.2. Causal Dependencies and DAGs
Let’s go back to the example of two observed variables $X$ and $Y$, which we briefly introduced in Section 1. Imagine that the process that determines the values of $X$ and $Y$ is perfectly known. In turn, we consider each of the three processes described by the equations in Figure 1. For simplicity, all our examples use linear dependencies and normal distributions, but the same line of thought is applicable to more complex, non-linear dependencies.
Process $p_{1}$ in 1(a) produces values of $X$ that are drawn randomly from a normal distribution with zero mean and unit standard deviation; and values of $Y$ that are perfectly proportional to those of $X$. Even in such an ideal scenario, any empirical measure of $X$ and $Y$ will include some measurement error $\epsilon$, which process $p_{1}$ models as another normal random variable that, together with $X$, determines the value of $Y$. We stress that we interpret 1(a)’s equations as capturing the causal dependencies among variables: $X$ and $\epsilon$ are drawn randomly (and independent of each other), and their random values determine (“cause”) the value of $Y$ in each draw. Correspondingly, the DAG (directed acyclic graph [36]) in 1(a) captures this causal relation between $X$ and $Y$ in a qualitative way: an edge connects $X$ to $Y$ to denote that the values of $X$ and $Y$ are related; furthermore, the edge is directed from $X$ to $Y$ to denote that changing $X$ directly affects $Y$, that is, it causes $Y$ to change.
Process $p_{2}$ in 1(b) involves a third variable $Z$. Just like variable $X$, $Z$’s values are drawn randomly from a normal distribution with zero mean and unit standard deviation. Then, the values of $Y$ are perfectly proportional to a linear combination of $X$ and ZZ—still with a term $\epsilon$ to account for measurement errors. Variables $X$ and $Z$ are independent of each other; the DAG in 1(b) clearly shows this independence, since it does not include any edge between $X$ and $Z$.
Process $p_{3}$ in 1(c) still involves the three variable $X$, $Y$, and $Z$. Now, $Z$ is the only variable that is drawn independently; in contrast, $X$ depends linearly on $Z$, and $Y$ depends linearly on a combination of $X$ and $Z$. As usual, the DAG in 1(c) visualizes these relations among variables, showing, in particular, that there is both a direct relation between $X$ and $Y$ (edge $X\to Y$) and an indirect relation through $Z$ (path $\underrightarrow{X\leftarrow Z\to Y}$).
2.3. Inference and Confounders
In an empirical study, the goal is to estimate, from a sample of the data, the parameters of a statistical model that captures the relations among observed variables. Let’s shift perspective on our illustrative examples: now, we are given a data sample $D_{1}$, $D_{2}$, $D_{3}$ respectively produced by each process $p_{1}$, $p_{2}$, $p_{3}$. Each data sample consists of many triples $(x_{i},y_{i},z_{i})$ of concrete values taken by $X$, $Y$, and ZZ [^2] Process $p_{1}$’s data sample only consists of pairs, since this process does not include a variable $Z$. in the $i$th observation—that is, the process’s $i$th draw. Now, we are imagining that we do not know the equations governing the generation process, but we want to quantitatively estimate the relation between $X$ and $Y$ from the observed sample.
Given that we are dealing with linear relations and normal distributions, we will use a regression model to fit the data. A key choice is whether to include variable $Z$ in the model: regression model $m_{1}$, shown in 2(a), ignores $Z$, whereas model $m_{2}$, shown in 2(b), includes $Z$ as a predictor. In both cases, after fitting a model on the data, the value of parameter $\beta$ will be the model’s estimate of the corresponding $b$ in Figure 1—in other words, $\beta$ estimates the causal “effect” of the predictor $X$ on the outcome $Y$. Table 1 shows the result of this exercise, when Figure 1’s processes use concrete values $b=0.4$, $c=0.7$, and $d=0.2$ for their parameters.
process
model $p_{1}$ $p_{2}$ $p_{3}$
$m_{1}$
$m_{2}$
(a) Whether each model $m$ estimates correctly () or incorrectly () the effect $b$ of $X$ on $Y$ in process $p$.
$\beta$ $\gamma$ $\sigma$
model $p_{1}$ $p_{2}$ $p_{3}$ $p_{1}$ $p_{2}$ $p_{3}$ $p_{1}$ $p_{2}$ $p_{3}$
$m_{1}$ 0.40 0.40 0.54 1.00 1.22 1.22
$m_{2}$ 0.40 0.40 0.70 0.70 1.00 1.00
(b) For each process $p_{1}$, $p_{2}$, $p_{3}$ the values of parameters $\beta$, $\gamma$, $\sigma$ in model $m_{1}$ or $m_{2}$ fitted on data generated by the process. In these experiments, the parameters of Figure 1’s processes are set to $b=0.4$, $c=0.7$, $d=0.2$.
Table 1. Estimating the parameters of Figure 1’s processes $p_{1}$, $p_{2}$, and $p_{3}$ with Figure 2’s regression models $m_{1}$ and $m_{2}$.
Process $p_{1}$.
The case of process $p_{1}$ is unproblematic: since the process does not involve any variable other than $X$ and $Y$, regression model $m_{1}$ accurately infers the value of parameter $\beta\simeq 0.40=b$, reflecting the true dependency between $X$ and $Y$. The fitted regression model also accurately infers the standard deviation $\sigma=1.00$ of the error term $\epsilon$. Obviously, model $m_{2}$ is inapplicable to analyze data produced by $p_{1}$, since this includes no variable $Z$.
Process $p_{2}$.
The case of process $p_{2}$ is more interesting, since we may analyze its data using either model $m_{1}$ or model $m_{2}$. As we expect, regression model $m_{2}$ accurately infers the value of parameter $\beta=b=0.40$, again reflecting the true dependency between $X$ and $Y$. Somewhat less obviously, the simpler model $m_{1}$ still infers the same correct value of parameter $\beta=b=0.40$, even if $Y$ also depends on $Z$ in generation process $p_{2}$. With model $m_{1}$, the effect of $Z$ on $Y$ has spilled into the estimate of the standard deviation $\sigma$ of the error term, which is in fact equal to $1.22$, greater than the “true” error standard deviation 1.01.0.333Equivalently, we can rewrite 1(b)’s generative equations as $Y=bX+E$, where $E\sim\mathrm{Normal}(0,\sqrt{1+c^{2}})$; if $c=0.7$, $\sqrt{1+c^{2}}\simeq 1.22$, which is exactly the inferred value of $\sigma$ in $m_{1}$ fitted on data from $p_{2}$.
Process $p_{3}$.
Regression model $m_{1}$ cannot accurately account for the more intricate data dependencies of process $p_{3}$. In fact, it overestimates the effect $\beta=0.54>0.40=b$ of $X$ on $Y$; now, this effect also includes the “spurious” correlation introduced by $Z$, which simultaneously affects $X$ and YY—as 1(c)’s DAG clearly shows. Variable $Z$ is called a confounder, since it mixes up the true effect of $X$ on $Y$ (path $X\to Y$ in the DAG) with an indirect, spurious correlation (path $\underrightarrow{X\leftarrow Z\to Y}$ in the DAG) that does not correspond to an actual data dependency but is just a figment of using an inadequate statistical model. Model $m_{2}$ makes up for $m_{1}$’s shortcomings by including $Z$ among its predictors; this is enough to cancel $Z$’s confounding effect on the link between $X$ and $Y$. Indeed, 1(b) shows that all parameters—crucially, the effect $\beta=0.40=b$ of $X$ on YY—are correctly estimated.
Using a model such as $m_{1}$ to analyze data from a process whose dependencies include a confounder (like process $p_{3}$) is an instance of omitted variable bias. The rest of the paper describes more realistic case studies where omitted variable bias may occur, and presents various mitigation strategies to counter the bias and recover a precise estimate of the causal effect linking treatment $X$ and outcome $Y$.
From the toy examples of this section, we can start to glean how omitted variable bias is commonplace in realistic settings. Even when our empirical data are rich and include many different variables, there is always a chance that we are missing some other variables that, like $Z$, confounds the effect of interest. Even if we are aware of possible confounders, measuring them to include them in our model (like model $m_{2}$ does) may be expensive, impractical, or impossible. For example, the confounder may lack a good operationalization, or it may be inaccessible because its values were not recorded and the process is not repeatable. These observations motivate the main contributions of the paper, which demonstrate how to identify and mitigate the wicked effects of confounders in a variety of practical scenarios.
2.4. Other Forms of Confounding
The term “confounding” is used to denote different, related concepts in the statistical analysis of empirical data [17].
  • In this paper, we use “confounding” to denote bias in the estimate of a causal effect—as we demonstrated in a nutshell in the previous sections. This notion of confounding is customary in modern causal analysis [36], and in related approaches to mitigate confounding, such as instrumental variables [2].
  • An early usage of “confounding” in statistics [49] was to denote noncollapsibility: roughly speaking, the association (correlation) between two variables is noncollapsible if its magnitude depends on whether we condition or not on another variable. Thus, noncollapsibility may denote a spurious causal effect [53], as in the previous meaning of “confounding”, but may also arise with different kinds of relations among variables; conversely, causal confounding may or may not determine noncollapsibility [17].
  • In classical frequentist statistics, the term “confounding” often denotes scenarios where main effects and interactions are inseparable [13]. If the interactions are not of interest, a study may deliberately introduce this kind of confounding, so that the main inferred effects summarize true effects and interactions.
  • In fields such as psychometrics, confounding is often characterized as a measurement problem or, more generally, an issue of experimental design [9]. In the present paper, in contrast, we take the main point of view of analyzing observational data, having little or no control on the process that generated the data.
While all these senses of the term “confounding” are related, the causal meaning that we follow in this paper captures a key challenge the analysis of observational data, and provides the clearest characterization of the omitted variable bias.
3. Confounding in Programming Languages and Code Quality
Our first case study follows closely the fundamental structure of Section 2.3’s prototypical example, while recasting it in a more realistic setting. Our goal is estimating the effect of using different programming languages (predictor variable Language) on the quality of the produced code written in that language (outcome variable Quality).
LanguageQualitySkill
Figure 3. The effect of Language on code Quality is confounded by the programmer’s Skill.
As we discussed in depth in related work [14], [^4] In this section, we analyze a subset of the same data collected by others [38, 4] that we also analyzed in our previous work [14]. This section’s and our previous work’s analysis [14] are, however, orthogonal: in this section, we analyze possible confounding effects based on causal assumptions; in contrast, [14]’s analysis is purely statistical and targets Bayesian statistical modeling practices. a programmer’s ability (expressed by variable Skill) is likely to confound the causal effect of Language on Quality. [^5] It’s plausible several other confounders of this causal relation exist [14]; for clarity of presentation, we only consider Skills, as if it captures the effects of other possible confounders of the same relation. Namely, a more skilled programmer is likely to produce higher-quality code ($\textnormal{{Skill}}\to\textnormal{{Quality}}$); and programmers with different abilities may prefer to work with certain programming languages over others ($\textnormal{{Skill}}\to\textnormal{{Language}}$). Figure 3 summarizes these relations by means of a DAG, which is isomorphic to 1(c)’s abstract DAG.
As discussed in Section 2.3, this DAG structure entails that, if we want to estimate the true effect of Language on Quality for observational data—where we cannot control the effect of Skill on Language, that is, we cannot randomize which language each programmer will use—we need to also condition on the confounder Skill. In practice, this may be impossible because precisely measuring Skill is not easy: for example, if the data comes from a source code repository, the identity of the programmers may be unknown; even if we have access to a programmer’s identity, it may be practically cumbersome to reliably assess their programming skills.
The rest of this section demonstrates how we can mitigate the effects of this unobserved variable bias, in a way that we can still get something out of our observational data about languages and code quality—even in such a tainted scenario.
3.1. Data: Programming Languages and Code Quality
To get a plausible ground truth about the relative effects of skills and languages on code quality, we analyze a subset of the data collected for a large-scale repository study [38], as made available in Berger et al. [4]’s reanalysis. This dataset summarizes the commit history of hundreds of projects in various languages.
To keep things simple—and to avoid bias that may come from underrepresented or misclassified languages [4]—we only retain data about projects: (1) written in Python or Java (two widely used, yet fairly different languages); (2) with at least 100 commits on record; (3) that are not multi-language (that is, each project is entirely written in Java, or entirely written in Python). These criteria select 105 projects: 45 written in Java and 60 written in Python. For each of these projects, we are interested in two key variables:
Language::
a binary, ordinal variable (values: Java or Python), which denotes the project’s language;
Quality::
a continuous variable (ranging over $[0,1]$), which measures the project’s quality as the complement $1-\textnormal{{Bugs}}/\textnormal{{Commits}}$ of the fraction of all project commits that are flagged as introducing a bug.
This dataset also includes information about which developer produced each commit. We use it to derive a crude estimate of the skills of developers active on a project as follows. First, we only consider the 952 developers who produced at least 10 commits each in the selected projects. The skill of each developer $d$ among these “frequent committers” is the complement $1-\textnormal{{Bugs}}_{d}/\textnormal{{Commits}}_{d}$ of the fraction of all commits authored by $d$ that introduced a bug. Finally, variable Skill summarizes the skills associated with each project:
Skill::
a continuous variable (ranging over $[0,1]$), which is the mean skill of all developers among the “frequent committers” who contributed to the project.
We stress that the goal of this data selection process is not supporting any general claims about the actual effects of a programming language on a project’s quality. It simply gives a rough idea of the magnitude of these effects in a real world scenario, so that we can appreciate that confounding is a plausible occurrence—hence, a practical concern.
3.2. Quantifying Confounding
In this exercise, the goal is estimating the effect of choosing Java or Python as a programming language (variable Language) on the quality of the developed project (variable Quality). If we had access also to variable Skill, we could single out the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect by fitting 4(a)’s regression model $m_{3}$, which uses Language and Skill as predictors. On the data described in Section 3.1, this produces an estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ of $\ell=-0.012$, which indicates that using Python is very weakly associated with a modest reduction of quality (with a lot of uncertainty, as shown in 2(a)).
$\textnormal{{Quality}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+\ell\cdot\textnormal{{Language}}_{i}+s\cdot\textnormal{{Skill}}_{i}$
(a) Model $m_{3}$ with two predictors Language and Skill.
$\textnormal{{Quality}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+\ell\cdot\textnormal{{Language}}_{i}$
(b) Model $m_{4}$ with one predictor Language.
$\textnormal{{Quality}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+s\cdot\textnormal{{Skill}}_{i}$
(c) Model $m_{5}$ with one predictor Skill.
model estimate 95% prob. 60% prob.
$m_{3}$ uncounfounded -0.012 -0.035 0.011 -0.021 -0.002
$m_{4}$ confounded -0.052 -0.094 -0.010 -0.070 -0.034
(a) Estimates of regression coefficient $\ell$ in models $m_{3}$ and $m_{4}$, with 95% and 60% probability intervals.
Figure 4. Different models to estimate the dependence between Quality and Language.
In practice, it may not be possible to reliably measure the confounder Skill. In this case, we can only fit 4(b)’s model $m_{4}$, which gives us a different estimate $\ell=-0.052$ of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ of Language on Quality. If we compare this with the previous estimate based on the unbiased model $m_{3}$, we notice that the confounding of Skill inflates the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ as measured in this data. In reality, we would not have access to the unbiased estimate; how to assess how much confidence we can put in an estimate that comes from a possibly confounded model? There are three main ways of proceeding [27]:
  • If we can muster an estimate the confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$, we can use it to compute how strong the effect $\textnormal{{Skill}}\to\textnormal{{Language}}$ would have to be to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect (i.e., flip it from negative to positive). Section 3.2.2 discusses this scenario.
  • Conversely, if we can estimate the confounding effect $\textnormal{{Skill}}\to\textnormal{{Language}}$, we can compute how strong the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ would have to be to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. Section 3.2.3 discusses this scenario.
  • Finally, assuming that there are several different confounders that simultaneously affect Language and Quality, we can compute how many confounders with a certain effect would it take to tip the estimate of the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. Section 3.2.4 discusses this scenario.
Assessing when tipping may occur under plausible scenarios can inform us about the plausible practical impact of the confounder on our data analysis: if tipping requires a strong confounding effect, which is unlikely in practice, we can probably tolerate the bias and still use our estimate, albeit imperfect, of the effect of Language on Quality. Conversely, if even a weak confounder is likely to tip, we would conclude that we should not rely on the observational data analysis, and try to collect different kinds of data (or to use the existing data to answer different kinds of questions). Such an analysis is useful also in more open-ended scenarios, for example when we don’t really know what (other) factors may affect Language and Quality but we have reasons to believe that some unmeasured variables exist.
To perform these so-called tipping point sensitivity analyses we will use the R package tipr [32], which implements state-of-the-art analysis techniques [27] based on causal models similar to those outlined in Section 2.2.
3.2.1. Scaled-mean Difference
In a tipping-point sensitivity analysis [32], any unmeasured confounder $Z$ is modeled as a standardized, normally distributed random variable. For a binary treatment variable $X$ with two nominal values $X_{0},X_{1}$, this means that the impact of the confounder $Z$ on the treatment $X$ is captured by normal distributions with unit variance and mean $\mu_{0}$ (when $X=X_{0}$) and $\mu_{1}$ (when $X=X_{1}$).
Accordingly, we use the scaled-mean difference (scd) to quantifying the impact of a confounder (Skill in our example) on the treatment (Language in our example) in a sensitivity analysis [32]. [^6] The scaled-mean difference $\mu_{1}/\sigma_{1}-\mu_{2}/\sigma_{2}$ between two groups is similar to, but different from, the standardized mean difference $(\mu_{1}-\mu_{2})/\sigma$—a commonly used effect size. Precisely, the scaled-mean difference is simply the difference of standardized means in the treatment and control groups. In our case, the $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})$ of a programmer’s skills in different language is the difference between the mean Skill of a Python programmer and the mean Skill of a Java programmer, expressed as a multiple of Skill’s standard deviation:
(1) $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})\quad=\quad\frac{\mathds{E}[\textnormal{{Skill}}\mid\text{Python}]}{\sigma[\textnormal{{Skill}}\mid\text{Python}]}-\frac{\mathds{E}[\textnormal{{Skill}}\mid\text{Java}]}{\sigma[\textnormal{{Skill}}\mid\text{Java}]}$
In (1), $\textnormal{{Skill}}\mid x$ denotes the values of Skill in all datapoints where $\textnormal{{Language}}=x$, whereas $\mathds{E}[D]$ denotes the mean and $\sigma[D]$ the standard deviation of some data $D$.
Since it’s based on a standardized scale, the scd is easy to interpret uniformly, across different settings and domains. For instance, an scd of $\pm 3$ denotes a huge effect: three standard deviations of difference indicate that the distribution of skills of Python and Java programmers barely overlap; clearly, such a massive difference in skills is unlikely to happen in practice. In our dataset, $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})$ is 1.545-1.545, which denotes quite a sizeable, albeit not huge, effect size.
measured estimate tipping
effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ -0.052 effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ 0.835 scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ -0.062
effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ -0.052 scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ -1.545 effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ 0.034
effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ -0.052 effect $\textnormal{{C}}\to\textnormal{{Quality}}$ 0.170 number of confounders C 2
scd $\textnormal{{C}}\to\textnormal{{Language}}$ -0.150
Table 2. Three scenarios where we calculate the effect sufficient to produce tipping the measured effect $\textnormal{{Language}}\to\textnormal{{Quality}}$, based on an estimate of the confounder Skill on either the outcome (first row) or the treatment (second row), or of several unknown confounders C (bottom rows).
3.2.2. Confounding scd
In this scenario, we have measured the (possibly confounded) effect $\textnormal{{Language}}\to\textnormal{{Quality}}$, and we have an idea (for instance, from other studies) of a plausible value for the confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$. From this data, we calculate the $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})$ that would lead to a confounding such that our estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ has the opposite sign of the “true” effect.
In our running example, the estimate of $\ell=-0.052$ using model $m_{4}$ is the measured effect $\textnormal{{Language}}\to\textnormal{{Quality}}$; whereas the estimate of $s=0.835$ using 4(c)’s model $m_{5}$ gives us a plausible value for the confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$. As shown in Table 2, a modest $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})$ of 0.062-0.062 would be sufficient to flip the sign of the measured effect. Such an scd is fairly modest, and likely to happen in practice; in fact, we have seen that the scd measured in the data is much larger. In all, we cannot have much confidence that the estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ is valid.
3.2.3. Confounding Effect
In this scenario, we have measured the (possibly confounded) effect $\textnormal{{Language}}\to\textnormal{{Quality}}$, and we have an idea (for instance, from other studies) of a plausible value for the confounding scd $\textnormal{{Skill}}\to\textnormal{{Language}}$. From this data, we calculate the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ that would lead to a confounding such that our estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ has the opposite sign of the “true” effect.
In our running example, the estimate of $\ell=-0.052$ using model $m_{4}$ is, once again, the measured effect $\textnormal{{Language}}\to\textnormal{{Quality}}$; whereas the $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})=-1.545$ measured on the data according to (1) gives us a plausible value for the confounding effect $\textnormal{{Skill}}\to\textnormal{{Language}}$. As shown in Table 2, a modest effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ of $0.034$ would be sufficient to flip the sign of the measured effect. Such a confounding effect is fairly modest, and likely to happen in practice; in fact, we have seen that the estimate of this effect $s=0.835$ using 4(c)’s model $m_{5}$ is much larger. Also in this scenario, we cannot have much confidence that the estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ is valid.
3.2.4. Number of Confounders
In this scenario, we have measured the (possibly confounded) effect $\textnormal{{Language}}\to\textnormal{{Quality}}$, and we are considering several different confounders. We have an idea of a plausible value for both the scd $\textnormal{{C}}\to\textnormal{{Language}}$ and the effect $\textnormal{{C}}\to\textnormal{{Quality}}$ for any such unknown confounders C. From this data, we calculate how many such variables C would produce an overall confounding such that our estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ has the opposite sign of the “true” effect.
As usual, the estimate of $\ell=-0.052$ using model $m_{4}$ serves as the measured effect $\textnormal{{Language}}\to\textnormal{{Quality}}$. Then, we can speculate that a generic confounder C has an $\textnormal{{scd}}(\textnormal{{C}}\to\textnormal{{Language}})=-0.15$ and an effect $\textnormal{{C}}\to\textnormal{{Quality}}=0.17$. These values are respectively $1/10$ and $1/5$ of the corresponding values for Skill as measured in the dataset; intuitively, they represent confounders with a much more tamed power compared to Skill. Nevertheless, just two such generic confounders would be sufficient to flip the sign of the measured effect. Since it is definitely plausible that there are a couple of confounders with moderate effect, we reach again the same conclusion that we cannot have much confidence that the estimate of the effect $\textnormal{{Language}}\to\textnormal{{Quality}}$ is valid.
Figure 5. A graphical summary of tipping analyses of confounder effect of Skill on the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. For different measured values of effect ($\beta_{\textnormal{{Language}}\to\textnormal{{Quality}}}$), the plot shows the values of the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ ($\beta_{\textnormal{{Skill}}\to\textnormal{{Quality}}}$) and of the scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ ($\textsf{smd}_{\textnormal{{Skill}}\to\textnormal{{Language}}}$) that would tip the main effect. The two blue cross marks correspond to the two scenarios discussed in Section 3.2.2 and Section 3.2.3.
3.3. Sensitivity Analysis
Let’s generalize the tipping analysis described in this section beyond Ray et al. [38]’s data. First of all, consider a range of possible measured effects $\textnormal{{Language}}\to\textnormal{{Quality}}$, including both positive and negative values. For each of them, Figure 5 plots the combination of values of confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ and scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ that would tip the measured effect.
Naturally, since both confounding factors concur to introduce tipping, the more pronounced one of them is, the less it is required of the other. Conversely, if the measured effect is small, even moderate-magnitude confounders may introduce a strong bias.
Such a sensitivity analysis can provide a useful guide not only to analyze empirical data, but to plan new experiments to improve the validity of existing findings. For example, it would be interesting to collect reliable data about the relations $\textnormal{{Skill}}\to\textnormal{{Quality}}$ (what’s the impact of a developer’s skills on the quality of code they produce?) and $\textnormal{{Skill}}\to\textnormal{{Language}}$ (do developers with different skill have marked preferences for which language to use?). Furthermore, a sensitivity analysis would enhance the value of an empirical study for other researchers, since it would better, and quantitatively, identify the study’s envelope of validity, and it would make the study’s assumptions more transparent.
4. Confounding in Teamwork Effort
Our second case study is based on Feldt et al. [11]’s review of research in the factors that affect the productivity of software development teams. It demonstrates how to practically assess the impact of potential unknown confounders in a more complex, realistic scenario. As in Section 3.2’s case study, our aim is not to (directly) contribute to the knowledge about developer productivity; in fact, our analysis will at times adopt simplifying assumptions that are not (entirely) realistic. In contrast, the goal is to demonstrate credible confounding patterns, and to illustrate how to navigate around them in realistic conditions.
STKDBEHLPO
Figure 6. A DAG summarizing relations among variables that characterize a software project, based on Feldt et al. [11]’s literature review. The variables are, from top to bottom and left to right: software Size, Team size, software Kind, Domain, Business area, Effort, Hardware, Location, Programming language, and Organization type.
4.1. Data: Teamwork, Effort, and Other Covariates
Feldt et al. [11] propose to use causal DAGs to summarize and combine the key findings of systematic literature reviews. In one of their case studies, they review several primary studies about productivity in software development. The DAG shown in Figure 6 is one of the DAGs that is obtained by applying Feldt et al. [11]’s approach; it summarizes the main relevant relations between variables that have been observed in some of the reviewed literature on the topic. [^7] This does not mean that this is the ultimate summary of research in the area of software development productivity. For our purposes, all that matters is that it displays a rich collection of plausible relations, so that our omitted variable bias analysis is grounded in a realistic scenario.
Figure 6’s DAG consists of 18 relations (arrows) among 10 variables:
B::
the company’s business area (accounting, sales, human resources, …)
D::
the project’s domain (healthcare, finance, entertainment, …)
E::
the overall effort spent by the software developers (hours worked, planned years, …)
H::
the hardware that runs the software (server, client, mobile, embedded, …)
K::
the kind of software (application, library, system, web, …)
L::
the company’s location (North America, Europe, Asia, …)
O::
the company’s organization type (private, public, non-profit, …)
P::
the project’s programming language
S::
the software size (lines of code, function points, …)
T::
the size of the team of programmers on the project
In the rest of this section, we put ourselves in the shoes of a researcher who is designing a new study to determine the strength of the causal relation between T (team size) and E (effort). The key questions that need to be addressed to design such a study are:
  1. i)
    What variables, other than the treatment T and the outcome E, ought to be measured?
  2. ii)
    Given the variables that could be effectively measured, what possible remaining confounders of the causal effect of treatment T on outcome E may remain?
Since Figure 6 summarizes several primary studies in the domain of software development productivity, we can conveniently use it as the basis of further studies in the same domain. Even if we were targeting a domain with little prior research, we could still build a DAG that captures whatever is known based on the state of the art in this domain. A DAG is just a convenient notation to summarize knowledge about the structural relations among variables; if previous work is scarce, the DAG will be simplistic or incomplete but will still serve as a useful guide. At a minimum, we can always fall back to building a minimal DAG such as in 1(c), which just captures the relation of interest $X\to Y$ and a generic confounder $Z$.
4.2. Adjustment Sets
As in the previous case study, to estimate the effect of a treatment (T in our example) on an outcome (E in our example), we fit on the data a linear regression model $m_{A}$:
(2) $\textnormal{{E}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+\beta_{t}\cdot\textnormal{{T}}+\sum_{v\in A}\beta_{v}\cdot v_{i}$
The estimate of coefficient $\beta_{t}$ in the fitted model $m_{A}$ measures the effect $\textnormal{{T}}\to\textnormal{{E}}$.
Thus, addressing question i) above (What variables ought to be measured?) is tantamount to deciding which variables (covariates) [^8] In common statistical jargon, a covariate is any predictor variable other than the treatment. should be included in set $A$ in model $m_{A}$ (2). In particular, $A$ should include all confounders, so that the effect $\textnormal{{T}}\to\textnormal{{E}}$ can be estimated without bias. Set $A$ represents what is called an adjustment set: given a DAG and two of its variables—representing the outcome (E in our case) and the treatment (T in our case)—the adjustment set $A$ is the set of additional variables in the DAG that we have to include in (2)’s model to ensure that $\beta$ estimates the unbiased, uncounfounded, genuine causal effect of T on E. In other words, the adjustment set includes all predictors that we should include if we want to avoid introducing omitted variable bias.
The adjustment set (or rather adjustment sets, since a DAG may admit multiple, alternative adjustment sets) can be computed directly on the DAG based solely on its structure—assuming, of course, that the DAG correctly captures real-world causal relations. [^9] In previous work [15], we illustrated how adjustment sets are computed, and what’s the intuition behind them. For brevity, we do not repeat the explanation in this paper, but simply use the adjustment sets computed from a DAG in our analysis. In our example, the adjustment set is $A=\{\textnormal{{O}},\textnormal{{S}}\}$; therefore including the two predictors O and S, in addition to T, ensures that $\beta$ captures the net causal relation between treatment and outcome.
4.3. Unmeasured Confounders
Even though Figure 6’s DAG is based on several empirical studies, it still is completely plausible that it does not include all factors that contribute to the observed relation between treatment and outcome. In fact, in every complex, real-world process, it is exceedingly likely that there are unmeasured variables that might still have a sizeable impact on the variables of interest.
confounded edge adjustment sets
1 $B\to E$ $\{O,S\}$
2 $B\to H$ $\{O,S\}$
3 $B\to P$ $\{O,S\}$
4 $B\to S$ $\{O,S\}$
5 $D\to E$ $\{O,S\}$
6 $H\to E$ $\{O,S\}$
7 $H\to P$ $\{O,S\}$
8 $K\to E$ $\{O,S\}$
9 $K\to S$ $\{O,S\}$
10 $L\to E$ $\{O,S\}$
11 $O\to E$ $\{O,S\}$
12 $O\to L$ $\{O,S\}$
13 $O\to P$ $\{O,S\}$
14 $O\to T$ $\{O,S\}$
15 $P\to E$ $\{O,S\}$
16 $S\to E$ $\{O,S\}$
17 $S\to T$ $\{B,K,O,S\}$ $\{O,S,Z_{S,T}\}$
18 $T\to E$ $\{O,S,Z_{T,E}\}$
Table 3. Adjustment sets for Figure 6’s DAG extended with a confounder $Z_{X,Y}$ affecting each edge $X\to Y$.
To address this issue of unmeasured (unknown) additional confounders, we can extend the previous adjustment set analysis. For every pair of nodes $X\to Y$ in Figure 6’s DAG, we introduce an unmeasured confounder $X\leftarrow Z_{X,Y}\to Y$ that affects $X$ and $Y$ simultaneously. Then, we recompute the adjustment set of the DAG extended with such additional node. Table 3 shows the results of this analysis.
Adding a confounder to 16 out of 18 edges in Figure 6’s DAG does not change the adjustment set, which remains $\{O,S\}$—as when considering the original DAG without unknown confounders. In other words, including $O$ and $S$ in model $m_{A}$ (2) conveniently also voids the effect of other possible confounders. In contrast, a confounder $Z_{S,T}$ affecting edge $S\to T$ admits two adjustment sets, one that includes $Z_{S,T}$ itself, and one that does not. Clearly, we prefer the latter adjustment set $\{B,K,O,S\}$, which enables us to correct the confounding effect of unknown variable $Z_{S,T}$ without even measuring it—in fact, without even knowing what this variable represents.
Unfortunately, the last edge $T\to E$ cannot be handled so easily. If there were an additional variable $Z_{T,E}$ simultaneously affecting $T$ and $E$, the only way to correct its confounding effect would be to measure $Z_{T,E}$ and include it as a predictor in (2)’s model. By definition, this is impossible because we don’t know what $Z_{T,E}$ is, or we have an idea but it is impractical or impossible to measure it.
[…]
Figure 7.  Figure 6’s DAG simplified to include only treatment $T$, outcome $E$, as well as all variables in the adjustment set that also accounts for unknown/unmeasured confounders on other edges. The DAG also shows an unmeasured confounder $Z$ that may still exist: we cannot adjust it away with any other variables.
4.4. Sensitivity Analysis
Figure 7 summarizes the analysis results so far: in order to get an unbiased estimate of the $T\to E$ causal relation in Figure 6’s DAG, we should include $B,K,O,S$ as additional predictors, which safeguards against unmeasured confounders affecting other relations in the DAG. However, there remains a possible confounder $Z_{T,E}$—which we’ll just call $Z$ from now on—that cannot be controlled for indirectly by means of other known variables.
Whether $Z$’s confounding is negligible or consequential ultimately depends on its strength relative to the strength of the other causal relations. Roughly, if $E\to T$ and the other relations are very strong, whereas $Z\to E$ and $Z\to T$ are very weak, our estimate of the $E\to T$ relation has a good chance of remaining reliable even if we cannot measure the unknown $Z$. Our next analysis step is thus a sensitivity analysis based on simulation: using some plausible, informed estimates for the various effects in the DAG, we’ll try to recover the effect $E\to T$ without measuring $Z$, and we’ll see how far off this estimate is from the ground truth.
4.4.1. Simulation Parameters
For our simulation, we use the following generative model, which mirrors the structure of Figure 7’s DAG:
[…]
$S$ $\sim\textsf{Normal}(b_{s}B+k_{s}K,1)$ $T$ $\sim\textsf{Normal}(o_{t}O+s_{t}S+\gamma_{t}Z,1)$
$E$ $\sim\textsf{Normal}(b_{e}B+k_{e}K+o_{e}O+s_{e}S+t_{e}T+\gamma_{e}Z,1)$
In (3), every categorical variable is binary and follows a Bernoulli distribution with $0.5$ probability of drawing a $1$; the other variables are real-valued and follow a normal distribution with unit variance and mean that is given by a linear combination of the variables that directly affect it according to Figure 7’s model. This is admittedly a strongly simplified generative model, but it has the advantage that we can choose standardized effect sizes for its parameters, instead of having to rely on difficult-to-obtain estimates on a natural scale. Furthermore, (3) could be generalized (e.g., to include categorical variables with more than two possible values), or specialized according to domain-specific characteristics (e.g., to use truncated distributions that capture hard bounds of the value of some variables).
The generative model in (3) has 11 parameters, one for each edge in Figure 7’s DAG. Each parameter $x_{y}$ denotes the effect of $X$ on $Y$, corresponding to edge $X\to Y$. The exceptions are edges $Z\to E$ and $Z\to T$, whose effects we denote as $\gamma_{e}$ and $\gamma_{t}$ to single them out (as they are the part of the model that is completely unmeasured). In a concrete case study, the simulation would use parameters that reflect the system that is actually being observed, where the data comes from. In our case, we do not have a specific data collection process in mind. Alternatively, one could simply try out all parameter combinations that are remotely plausible. In our case, such an exhaustive analysis would be practically infeasible; for example, if each parameter can take 6 possible values, we would end up with $6^{11}$ parameter combinations—that is over 362 millions!
Instead, we go back to Feldt et al. [11] and use the statistics from some of the reviewed primary studies to get a ballpark estimate of the relevant effects. This trades off some generality for a manageable simulation time. Table 4 shows the range of parameter values that we used for our simulations.
Confounder::
First of all, we want to simulate all plausible confounding scenarios of $Z$; therefore, we consider all combinations of small (0.1), medium (0.3), and large (0.5) standardized effect sizes [^10] The values 0.1, 0.3, and 0.5 are often considered as the boundaries between negligible, small, medium, and large effect sizes [8, p. 224–225] [12]. —both positive and negative—for each parameter $\gamma_{e}$ and $\gamma_{t}$.
Main effect::
We also consider all possible effect sizes for parameter $t_{e}$, which represents the treatment effect that we are trying to estimate; however, we only consider positive effect sizes since, according to the studies reviewed in Feldt et al. [11], it is implausible that large teams produce an overall lesser effort than small teams.
Indirect effects::
As for the effects $x_{e}$ corresponding to the edges $X\to E$, for every other variable $X$ in the adjustment set, we picked a small, medium, or large, positive or negative effect size based on some of the studies reviewed in Feldt et al. [11]. Since those studies focused on effort (or related variables) as outcome, we could not find any hard data about the magnitude of the effects of these other variables $X$ on other covariates. Simplistically, we assume that $x_{y}$ is the same as $x_{e}$, that is variable $X$ has roughly the same effect on all variables it directly affects. Again, a specific case study could come up with more definite estimates; our goal is mainly to demonstrate this analysis method on somewhat plausible data.
parameter values justification
$b_{e}$ $B\to E$ 0.3 [48, Tab. V, $\omega^{2}$/maintenance (all)]
$b_{s}$ $B\to S$ 0.3 Same as $b_{e}$
$k_{e}$ $K\to E$ 0.1 [52, Tab. XIII, % variance explained/project type]
$k_{s}$ $K\to S$ 0.1 Same as $k_{e}$
$o_{e}$ $O\to E$ 0.5 [48, Tab. IV, $\omega^{2}$/maintenance (all)]
$o_{t}$ $O\to T$ 0.5 Same as $o_{e}$
$s_{e}$ $S\to E$ -0.1 [48, Tab. XIII, $\rho$/maintenance (all)]
$s_{t}$ $S\to T$ -0.1 Same as $s_{e}$
$t_{e}$ $T\to E$ 0.1, 0.3, 0.5 all plausible positive effect sizes
$\gamma_{e}$ $Z\to E$ -0.5, -0.3, -0.1, 0.1, 0.3, 0.5 all plausible effect sizes
$\gamma_{t}$ $Z\to T$ -0.5, -0.3, -0.1, 0.1, 0.3, 0.5 all plausible effect sizes
$n$ sample size 5, 10, 50
$n_{\textrm{sim}}$ repetitions 200
Table 4. Range of values for the parameters of the generative model in (3) used in the simulation. For each parameter, the table also reports the justification for the choice of values, usually as a reference to a primary study that measured such an effect or a comparable one.
The simulation includes implicitly two more parameters:
Sample size::
the sample size $n$ determines how many datapoints we sample from (3)’s generative model. We try three different sample sizes: 5, 10, and 50. Due to the nature of the data we are simulating, small sample sizes (i.e., 5 and 10) are especially relevant and realistic: collecting all such detailed data about many software projects would be costly (in particular, it’s unlikely that such data can be reliably obtained by simply mining open-source repositories); hence, an actual empirical study would likely be limited to a smallish sample size. Nevertheless, we also include a more substantial sample size (i.e., 50) to extend the reach of our analysis.
Repetitions::
For each parameter combination, we repeat the whole simulation-inference process $n_{\textrm{sim}}$ times, and take the average of the obtained estimates. We go with 200 repetitions, which should be enough to smoothen out any random fluctuations in our simulations.
4.4.2. Simulation Process
For every combination of values for the parameters in Table 4, Algorithm 1 uses generative model (3) to draw random samples of the variables $B,K,O,Z,S,T,E$ (consistently with the dependencies shown in Figure 7). It then fits the following linear regression model on the samples:
(4) $\textnormal{{E}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
$\mu_{i}$ $=\alpha+\beta_{t}\cdot\textnormal{{T}}_{i}+\beta_{b}\cdot\textnormal{{B}}_{i}+\beta_{k}\cdot\textnormal{{K}}_{i}+\beta_{o}\cdot\textnormal{{O}}_{i}+\beta_{s}\cdot\textnormal{{S}}_{i}$
Model (4) estimates the effect size $\beta_{t}$ of $T$ on $E$ without conditioning on $Z$, which we assume cannot be measured, but otherwise includes all measurable variables in the adjustment set.
This sample/fit process is repeated $n_{\textrm{sim}}$ times for each parameter combination; finally, Algorithm 1 returns the mean $\beta_{t}$, the 50% highest-probability density interval $\ell_{50}\ldots u_{50}$, and the 95% highest-probability density interval $\ell_{95}\ldots u_{95}$ of the estimates of $\beta_{t}$ over all repetitions.
Figure 8. A summary of the results of running Algorithm 1 for some parameter combinations in Table 4. Each plot reports, for each sample size 5, 10, 50, the ground truth $t_{e}$ , the mean estimate , the 50% probability interval (shifted left), and the 95% probability interval (shifted right) of $\beta_{t}$.
Input: parameters $b_{s},b_{s},k_{e},k_{s},o_{e},e_{t},s_{e},s_{t},t_{e},\gamma_{e},\gamma_{t},n,n_{\textrm{sim}}$
Output: estimate $\beta_{t}$; 50% interval $\ell_{50}\ldots u_{50}$; 95% interval $\ell_{95}\ldots u_{95}$
$\mathit{est}\leftarrow\emptyset$
// repeat $n_{\textrm{sim}}$ times
for $r\leftarrow 1\ldots n_{\textrm{sim}}$ do
 $\mathit{sim}\leftarrow\emptyset$
 // collect $n$ random samples from generative model (3)
 for $s\leftarrow 1\ldots n$ do
    $B,K,O\leftarrow$ samples from $\textsf{Binomial}(1,0.5)$
    $Z\leftarrow$ sample from $\textsf{Normal}(0,1)$
    $S\leftarrow$ sample from $\textsf{Normal}(b_{s}B+k_{s}K,1)$
    $T\leftarrow$ sample from $\textsf{Normal}(o_{t}O+s_{t}S+\gamma_{t}Z,1)$
    $E\leftarrow$ sample from $\textsf{Normal}(b_{e}B+k_{e}K+o_{e}O+s_{e}S+t_{e}T+\gamma_{e}Z,1)$
    // $\mathit{sim}[s,V]$ stores the value of variable $V$ in the $s$th sample
    $\mathit{sim}[s,B],\mathit{sim}[s,K],\mathit{sim}[s,O],\mathit{sim}[s,Z],\mathit{sim}[s,S],\mathit{sim}[s,T],\mathit{sim}[s,E]\leftarrow B,K,O,Z,S,T,E$
   end for
 $f\leftarrow$ fit((4), $\mathit{sim}$) // fit model (4) with data $\mathit{sim}$
 // $\mathit{est}[r]$ stores the fitted model’s estimate of $\beta_{t}$ in the $r$th repetition
 $\mathit{est}[r]\leftarrow$ estimate($\beta_{t},f$)
end for
// compute statistics over all $n_{\textrm{sim}}$ repetitions
[…]
return $\beta_{t},\ell_{50},u_{50},\ell_{95},u_{95}$
Algorithm 1 Simulation algorithm to assess the sensitivity of the estimate of $T\to E$’s effect on the confounding effect of $Z$ in Figure 7.
4.4.3. Simulation Results
Figure 8 summarizes the results of the simulation. The figure only includes the results for the following values for effect sizes $\gamma_{e},\gamma_{t}$: -0.5, 0.1, 0.3, 0.5. This subset of the results is sufficient to see the main trends in the sensitivity analysis; anyway, the replication package [16] includes plots for all parameter combinations. The capsule summary of these results is that it is possible to retrieve a reasonably precise estimate of the effect $t_{e}$ provided the confounding effect of the unmeasured $Z$ (parameters $\gamma_{e}$ and $\gamma_{t}$) is not large compared to $t_{e}$.
Let’s first look at the plot grid in the middle of Figure 8, which correspond to a “ground truth” medium effect $t_{e}=0.3$ (marks ). When the confounding effects are small ($\gamma_{e},\gamma_{t}=0.1$, second row or column of the grid) or medium ($\gamma_{e},\gamma_{t}=0.3$, third row or column of the grid), the estimates (marks ) are quite close to the actual effect, nearly overlapping it. Remarkably, this holds even for only 5 or 10 samples; however, a small sample size leads to very wide probability intervals, even for 50% probabilities (left of the estimate). Increasing the sample size to 50 substantially shrinks the probability intervals; if obtaining a substantial number of datapoints is challenging in practice, these results indicates that a substantial uncertainty about the accuracy of the estimate would remain. In contrast, as we consider larger confounding effects ($\gamma_{e},\gamma_{t}=\pm 0.5$, first and last row or column of the grid), there is a substantial gap between the estimates and the actual effect . Precisely, when $\gamma_{e}$ and $\gamma_{t}$ have the same sign (both positive, or both negative), $Z$’s bias results in overestimating the ground truth effect; conversely, when $\gamma_{e}$ and $\gamma_{t}$ have opposite sign, the bias results in underestimating the ground truth effect.
If we now consider a “ground truth” small effect $t_{e}=0.1$ (plot grids in the top of Figure 8) or a large effect $t_{e}=0.5$ (plot grids in the bottom of Figure 8) we largely see the same trends. On the one hand, a large effect size is not strictly “harder to bias”; that is, the estimate of $t_{e}=0.5$ is biased in a roughly similar way by certain confounding effects $\gamma_{e},\gamma_{t}$ as the estimate of $t_{e}=0.1$. Intuitively, this happens because we still condition on all other variables in the adjustment set, and hence what is left is the net confounding effect of omitting $Z$ over the estimate. On the other hand, if the “true” effect we are estimating is small compared to the biasing influence of $Z$, the same absolute amount of bias translates into an estimate error that is possibly much more consequential: even the 50% probability intervals clearly overlap zero; thus, in a real setting it would be hard to conclude that a definite (positive or negative) effect exists at all.
4.5. Sensitivity Analysis with E-values
The sensitivity analysis based on simulation presented in Section 4.4 is informative and flexible, but it also has clear disadvantages: it can be very time consuming, and it does usually requires a good amount of empirical data to tune the simulation parameters in a way that is representative of the domain. As a less demanding alternative, this section presents a tipping-point analysis similar to the one we described in Section 3.2 for the other case study.
4.5.1. E-values
Section 3.2’s sensitivity analysis uses the scd (scaled-mean difference) as a standardized measure of the effect of an unmeasured confounder on the treatment. The scd (1) is defined based on a dichotomous partition of the treatment variable. In Section 3’s domain, the treatment was the programming language, which is naturally dichotomous. In contrast, team size $T$ (the treatment variable in our current domain) is an intrinsically quantitative (numeric) variable; in order to calculate an scd of $Z\to T$, we would have to arbitrarily partition teams into small vs. large. Instead, we rely on a different kind of tipping-point analysis based on the notion of E-value [51], which is also applicable to continuous quantitative exposure variables. [^11] We rely on the R package EValue [29], which implements a variety of state-of-the-art sensitivity analysis techniques for unmeasured confounding based on the notion of E-value [6, 51, 50].
The E-value [^12] The “E” stands for “Evidence”. is “the minimum strength of association, on the risk ratio scale, that an unmeasured confounder would need to have with both the treatment and the outcome to fully explain away a specific treatment-outcome association, conditional on the measured covariates” [51]. In our scenario, an estimate $\beta_{t}$ of the observed effect $T\to E$, obtained by fitting model (4) on empirical data, would represent the “specific treatment-outcome association, conditional on the measured covariates.” The E-value can be computed from $\beta_{t}$, as well as $\epsilon_{t}$ (the standard error of $\beta_{t}$’s estimate), an estimate of $\sigma$ (parameter $\sigma$ in model (4)), and a parameter $\delta$ that quantifies the arbitrarily chosen change in treatment variable $T$. With these parameters, an E-value of $e$ can be interpreted as follows: if the unmeasured confounder $Z$ were strong enough to increase, by a factor of $e$, the probability of raising the exposure $T$ by $\delta$—while simultaneously affecting the outcome $E$ by a similar amount in standardized units—then the observed effect $\beta_{t}$ would be entirely due to $Z$’s confounding.
Values $\beta_{t}$, $\epsilon_{t}$, and $\sigma$ all come from fitting model (4) on empirical data; this makes the E-value a convenient way of estimating the effect of unmeasured confounders, since it is a byproduct of a standard regression analysis. Since it also depends on $\epsilon_{t}$ and $\sigma$, computing the E-value from a regression analysis brings the additional advantage that it takes into account the uncertainty in the estimates, as well as the other covariates the regression model conditions on. In contrast, parameter $\delta$ in the computation of the E-value can be chosen arbitrarily so that it reflects a variation “of interest” in $T$. Ultimately, $\delta$ still implicitly introduces a dichotomous partition of the treatment, but it does so in a way that is more apt for a continuous treatment.
4.5.2. Computing E-values
Let’s get into computing the E-value in our case study of team size and effort. Since we don’t have real-world data directly about the variables of Figure 7, we resort, once again, to simulation to get some plausible data in a convenient format. Unlike in Section 4.4, now we don’t need to perform many repetitions with small sample size, since the simulated data will not be used for a sensitivity analysis but to create a synthetic dataset based on the Table 4’s parameters. Thus, we simply draw ten thousand datapoints by sampling model (3) for each of the following parameter combinations:
  • $b_{e},b_{s},k_{e},k_{s},o_{e},o_{t},s_{e},s_{t},t_{e}$ are as in Table 4; namely, we consider a range of positive effect sizes for $t_{e}$, whereas we stick with realistic values for the other parameters.
  • $\gamma_{e}=\gamma_{t}=0$; in other words, we do not introduce any confounding in the simulated model. This simplifying assumption does not affect the following analysis, since it still makes sense to compute an E-value in such a scenario: the E-value quantifies the magnitude of a hypothetical confounder given an observed effect; it is not a way of detecting confounders but of reasoning about their possible strength. [^13] For completeness sake, the replication package includes a computation of E-values in scenarios where $\gamma_{e}$ and $\gamma_{t}$ are non-zero and range over the same values as in Table 4.
(a) Each colored line shows the E-value for each $\delta$ when the $T\to E$ effect is $t_{e}$.
$\delta$
$t_{e}$ 0.1 0.3 0.5
0.1 $1.10$ $1.19$ $1.26$
0.3 $1.20$ $1.39$ $1.55$
0.5 $1.26$ $1.54$ $1.80$
(b) E-values for different combinations of $T\to E$ effect $t_{e}$ and parameter $\delta$.
Figure 9. E-value tipping-point analysis for the unknown confounder $Z$ on the $T\to E$ relation in Figure 7.
Then, we fit model (4) with each of the three simulated datasets $D^{0.1},D^{0.3},D^{0.5}$ (one for each value of $t_{e}=0.1,0.3,0.5$). The fit with data $D^{s}$ gives values $\beta_{t}^{s}$, $\epsilon_{t}^{s}$, $\sigma^{s}$, respectively of the estimated effect $T\to E$, of the standard error of this estimate, and of the standard deviation of model (4)’s likelihood. With $\beta_{t}^{s}$, $\epsilon_{t}^{s}$, $\sigma^{s}$, we compute the E-value for many different values of parameter $\delta$ ranging from 0.01 to 0.5.
4.5.3. Confounding Sensitivity
9(a) plots the E-values for the three effect sizes $0.1$, $0.3$, and $0.5$. Overall, the E-value is proportional to $\delta$ (the parameter that captures the increase in treatment level caused by a possible confounder $Z$) and to $t_{e}$ (the actual effect $T\to E$). This simply reflects that the bigger the effect to be explained away by a confounder, the stronger the confounder has to sway treatment and outcome at the same time.
Then, 9(b) lists precise E-values for certain combinations of $\delta$ and $t_{e}$. Let’s look into a couple of interesting parameter combinations:
  1. i)
    When $t_{e}=\delta=0.1$, the E-value is $1.10$, which can be interpreted as follows: if $Z$ is such that, in nominal conditions, it can raise by at least 10% the probability of increasing a project’s team size by 0.10.1—while correspondingly increasing the effort by a “similar” amount—(where the increase is subject to some standard distributional assumptions), then $Z$ would be sufficient to explain entirely the observed effect $t_{e}=0.1$. In (3)’s standardized model, $0.1$ is a small increase; thus, it is not unrealistic that a certain unmeasured factor increases (by 10%) the probability of ending up with a moderately larger team size and effort.
  2. ii)
    When $t_{e}=0.5$ and $\delta=0.1$, the E-value is $1.26$. Informally, to explain away a large observed effect, $Z$ should increase by 26% the probability of introducing a small increase in team size ($\delta=0.1$). This is definitely a more significant impact than in the previous point i), but it is perhaps still plausible.
  3. iii)
    When $t_{e}=\delta=0.5$, the E-value is $1.80$. Informally, to explain away a large observed effect, $Z$ should increase by 80% the probability of introducing a large increase in team size. This scenario is no longer so plausible: it is a case of “extraordinary claims [requiring] extraordinary evidence” [41].
Comparing scenarios ii) and iii) is also interesting. Overall, they suggest that it is still possible that a confounder is responsible for a large effect; however, this is plausible only if the confounder $Z$ can bias the estimate of the $T\to E$ effect with only a small change in the treatment $T$ (scenario ii), where the $T\to E$ is more sensitive), as opposed to a large change (which is less likely to have been missed by the quantitative analysis).
As usual, with a better understanding of the domain (for example, the typical project characteristics in the company where the analyzed data was collected), the E-value analysis could produce more actionable results. For example, if one is confident that the estimates of team size and project effort that are normally produced are usually precise, it would indicate that it’s unlikely that a confounder would go undetected even for small changes of treatment/outcome (e.g., $\delta=0.1)$. Conversely, if the characteristics of the projects that are being analyzed make estimates intrinsically imprecise or uncertain, an unmeasured confounder that substantially increases the probability of swaying such estimates is a plausible possibility. As one example, Morasca and Russo [34] study of productivity (one of those surveyed in Feldt et al. [11]) presents estimates of productivity with a large dispersion ($\sigma\in[0.35,7.2]$); these indicate much uncertainty, thus raising the possibility of unmeasured confounders.
5. Dealing with Omitted Variable Bias
This section is a high-level summary of the techniques presented in the paper. The summary also serves as a procedural checklist, presenting the main steps of an analysis of confounding and the order in which they should be followed.
Variables.
The first step is surveying the variables that characterize the target of our study, their types (categorical, ordinal, numeric, …), how costly they are to measure (e.g., they can be mined from software repositories vs. they require running a controlled experiment), and how much uncertainty we expect in their measures. We should also select which variables are the treatment (a.k.a. exposure) and the outcome, whose relation is the main focus of the study. This step underlies all the following ones, as it provides a way of becoming familiar with the study’s domain in an incremental fashion.
Causal DAG.
The variables of interest, identified in the previous step, serve as nodes of a DAG such as those in Figure 3 and Figure 6. Causal DAGs are the notation of choice to succinctly express structural, causal relations among variables. As we have demonstrated in the paper, there are several analyses that are based on a DAG’s structure.
What kind of information we can use to build a DAG depends on the domain we are investigating, and on the maturity of the state of the art. If there are plenty of rigorous primary studies about the quantities of interest, and perhaps even systematic reviews or meta-studies, we can summarize their evidence in a DAG—similarly to what is proposed by Feldt et al. [11]. If our study’s target is novel or less established, we may have to rely on domain expertise and intuition to build a DAG. Even if there is a lot of uncertainty about the precise causal structure underlying a certain domain, there are techniques to (partially) validate candidate DAGs [15]. It is quite natural to also consider different possible DAGs, and to use them to perform a “what if” analysis in different scenarios. In this step of the analysis, causal DAGs are mainly a convenient notation to rigorously express our knowledge or hypotheses about the causal relations among variables, and to investigate their consequences on the overall results of our analysis.
Adjustment sets.
As shown in Section 4.2, given a DAG and treatment/outcome variables, one can systematically compute an adjustment set: a set of covariates that should be conditioned on in a regression to ensure that the coefficient associated with the treatment variable estimates the unbiased effect of the treatment on the outcome (“controlling for” the spurious influence of any confounders).
In the best-case scenarios, one can simply use an adjustment set to prevent confounding. Unfortunately, this is not always possible. First, the adjustment set’s validity is predicated on the accuracy and completeness of the DAG: if we missed some relevant variables, or misrepresented some causal relations, an adjustment set no longer guarantees an unbiased estimate. Second, even if we are confident the DAG is accurate, certain DAGs do not admit adjustment sets (because different kinds of confounding require incompatible adjustment sets [30]), or some variables in the adjustment set are hard or impossible to measure (a scenario that we explored in Section 4.3). In these cases, the next steps in this list can help deal with these shortcomings.
Ballpark estimate of parameters.
In order to proceed with a sensitivity analysis of unmeasured confounders, one needs to collect a rough estimate of the strength of the main relations among variables in the DAG. If a good amount of data is available from the study’s domain, we can use them to come up with estimates on the natural scale. Otherwise, we can still resort to mocking an ersatz model, based on the DAG’s relations, that uses standardized variables. On a standardized scale, it is easier to make “guesstimates” about plausible parameter values (e.g., small vs. large), and to explore the impact of different parameter combinations—as we did to select the parameter values in Table 4.
Sensitivity analysis.
Using the parameter estimates identified in the previous step, one can perform different kinds of sensitivity analyses of unmeasured confounder. These analyses provide a quantitative estimate (usually on a normalized scale) of how much some unmeasured confounder may bias the estimate of an effect of interest. In the paper, we demonstrated two kinds of so-called tipping-point sensitivity analysis, which express how strong an unmeasured confounder should be to cancel out an observed treatment/outcome effect. Section 3.2 presented a tipping-point analysis based on an estimate of the scaled-mean difference of an unmeasured confounder on the treatment, which is applicable to dichotomous treatment variables. Section 4.5 presented a sensitivity analysis based on E-values—a probability ratio between the confounded and non-confounded scenarios—which is also applicable to continuous treatment variables.
If a sensitivity analysis indicates that possible unmeasured confounders are unlikely to exist, or to have a noticeable impact, one can proceed with the real data analysis, reassured that confounding is a remote possibility.
Simulation analysis.
If the previous step’s sensitivity analysis is inconclusive, in that it failed to rule out the possibility of confounding, one can perform a more precise analysis of confounding based on simulation. Section 4.4 illustrated this on our second case study of teamwork productivity. A simulation analysis is flexible, because one can explore many different variants of generative and inference models. It also supports analyzing the impact of dealing with small sample sizes in a way that realistically reflects the availability of data in the study domain. These advantages come with a cost in terms of simulation time; usually, however, it still is much cheaper to perform a detailed simulation than to embark “blind” in running an empirical study without a clear understanding of the possible confounders, and of the threats to validity they may introduce.
Study design and execution.
All previous steps are ultimately a preparation for the design and actual execution of the envisioned study. Precisely, there are three main outcomes of the previous steps:
All clear::
the analysis indicates that confounding is not possible (because of the DAG structure), can be prevented (using a suitable adjustment set), or is unlikely to have a sizeable impact (as shown by the sensitivity analysis). This is the best-case scenario, which bodes well for the validity of our study.
Proceed with caution::
the analysis indicates that confounding is a possibility, but, depending on the effects that are in place and on the sample size that we may be able to collect, may or may not be consequential. In this case, we may still decide to go ahead with our study or, more cautiously, we may perform additional preliminary analyses (for example using detailed simulations) to gauge more precisely the quantitative relations that animate our domain.
No go::
the analysis indicates that major confounding is unavoidable, and that our estimates of effects are likely to be ridden with uncertainty. In this case, it may not be worth to proceed with the study as originally intended. Instead, we may refocus our goals, and redefine our research questions, so as to move them to a scope that is more likely to be productive.
6. Conclusions
This paper introduced to the empirical software engineering community techniques to assess and mitigate the so-called omitted variable bias. These techniques are crucially based on a causal model of the relations among variables of interest, formalized by means the causal DAG notation [36].
First, if the causal structure among variables admits an adjustment set that only includes measurable variables, one can correct for an omitted variable bias by simply conditioning on all variables in the adjustment set. If this is not possible, one can perform a sensitivity analysis, whose goal is investigating the impact of unknown confounders. The paper presented different kinds of sensitivity analyses, including both so-called tipping-point analyses based on canonical distributional assumptions, and more precise, but also computationally expensive, analyses based on simulation. We demonstrated these techniques on two case studies—the relation between programming languages and code quality, and the effect on team size on software development effort—taken from recent statistical analyses of data in these two domains [38, 11].
The main high-level takeaway of this work is to think before you act. The most effective way of designing an empirical study is to start with an elicitation of the causal model(s) that underlie the phenomena under study, followed by a systematic and explicit sensitivity analysis of possible confounders. This will lead to a clearer understanding of the limitations of a particular study and, in turn, to a more effective study design—one that is less likely to incur major threats to validity.